OpenStudy (shadydare):

Divide. 4/5 2/3 A. 8/15 B. 5/6 C. 1 1/5 D. 1 7/8 Please Help!

OpenStudy (flamingpiano):

I got C. 1 1/5

OpenStudy (anonymous):

@ShadyDare Are we dividing 4/5 by 2/3?

OpenStudy (shadydare):

yes

OpenStudy (mrnood):

if you are dividing by a fraction then that is the same as 'inverting' the fraction and then MULTIPLYING so \[\frac{ \frac{ 4 }{ 5 } }{ \frac{ 2 }{ 3 } } = \frac{ 4 }{ 5 } \times \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

C is correct. (12/15)/(10/15)=1.2=1 1/5

OpenStudy (shadydare):

okay thanks! :D

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }\div \frac{ 2 }{ 3 }\] We need to divide the first fraction by the second fraction. Step 1. We copy the first fraction as it is.\[\frac{ 4 }{ 5 }\]Step 2. We switch the division into a multiplication.\[\frac{ 4 }{ 5 }\times\]Step 3. We flip the second fraction.\[\frac{ 4 }{ 5 }\times \frac{ 3 }{ 2 }\]Step 4. We multiply the fractions.\[\frac{ 4 }{ 5 }\times \frac{ 3 }{ 2 }=\frac{ 4\times3 }{ 5\times2 }=\frac{ 12 }{ 10 }\]The multiplication ios done. Now we see if we can reduce the fraction. Step 5. Try to reduce the result. We notice that both 12 and 10 are divisible by 2.\[\frac{ 12 }{ 10 }=\frac{ 6\times2 }{ 5\times2 }=\frac{ 6 }{ 5 }\]

OpenStudy (shadydare):

wow thanks @MARC_

OpenStudy (anonymous):

yw! @ShadyDare